OpenStudy (anonymous):

Use and augmented matrix to solve the system. x + y = 5 3x - y = -1 From what I can I read, it seems it would be (3,5) but that is not one of the answer choices. I would rather have an explanation than an answer. I think the matrix is: [1 1]5] [3 -1]-1] Any help is greatly appreciated.

OpenStudy (fibonaccichick666):

That would be the matrix of coefficients

OpenStudy (fibonaccichick666):

and it is correct

OpenStudy (fibonaccichick666):

What you want to do is reduce the matrix using Reduced Row Echelon(I think thats the term or its Gauss but you get it) you want it of the form: \[\left[\begin{matrix}1 & 0 & x \\ 0 & 1 & y\end{matrix}\right]\]

OpenStudy (anonymous):

I found what to do in the back of the book, could you stick around and see if I get the right answer?

OpenStudy (fibonaccichick666):

sure

OpenStudy (fibonaccichick666):

Any progress?

OpenStudy (anonymous):

Well, apparently x=y+2.333.... I think I am lost lol

OpenStudy (fibonaccichick666):

yea, it's alright lets try from the beginning

OpenStudy (fibonaccichick666):

So you have :\[\left[\begin{matrix}1 & 1 & 5 \\ 3 & -1 & -1\end{matrix}\right]\] right?

OpenStudy (anonymous):

Yes. Now you multiply the second row by -2 and add it to the first right?

OpenStudy (fibonaccichick666):

to type the matrix, right click on mine and hit show math as tex commands it'll make your life easier. Why -2?

OpenStudy (fibonaccichick666):

and why second?

OpenStudy (anonymous):

That is what my book says...

OpenStudy (fibonaccichick666):

I disagree. Does your book have the answer too? ( I would like to check myself)

OpenStudy (anonymous):

It is not for this equation, it was the only thing I could find on matrices.

OpenStudy (fibonaccichick666):

oh ok not quite then. I was like I know I checked my answer at least 3 times... lol alright so in order to solve a matrix you want to get a 1 in the top left first. do we have that?

OpenStudy (anonymous):

Yes.

OpenStudy (fibonaccichick666):

now you want all zeroes directly below it, how do we achieve this? we have two weapons. 1. we can multiply the first row by something. 2. We can add the first row to any other row. So can you think of a way to get a zero where the three is?

OpenStudy (anonymous):

You can add -3 to get zero.

OpenStudy (fibonaccichick666):

right so what does that give you? remember you have to add the whole row

OpenStudy (anonymous):

That row would be [0 -4|-4]

OpenStudy (fibonaccichick666):

-4?

OpenStudy (anonymous):

-1+(-3)?

OpenStudy (fibonaccichick666):

remember the entire row, the columns stay together so the second column would be -4 but what about the third?

OpenStudy (anonymous):

Oh, doesn't that stay the same?

OpenStudy (fibonaccichick666):

nope same rule, as the others

OpenStudy (anonymous):

The 3rd would be -4 then?

OpenStudy (fibonaccichick666):

no, first what i 3*5?

OpenStudy (anonymous):

15

OpenStudy (fibonaccichick666):

k now what is -15+-1

OpenStudy (anonymous):

-16. Where did the 5 come from?

OpenStudy (fibonaccichick666):

The last column from the original matrix do you see it?

OpenStudy (anonymous):

That is the first row, arn't we working on the second though lol

OpenStudy (fibonaccichick666):

nope, remember if we are adding rows, we add the whole row so -3[ 1 1 5] + [3 -1 -1]

OpenStudy (anonymous):

Ohhhh, I get it now lol.

OpenStudy (fibonaccichick666):

good! yay! ;P

OpenStudy (fibonaccichick666):

so can you tell me what the second row becomes?

OpenStudy (anonymous):

[0 -4|-16]

OpenStudy (fibonaccichick666):

correct now can you make the first non-zero in the second row =1?

OpenStudy (anonymous):

Add 5 to it

OpenStudy (fibonaccichick666):

no remember you can only multiply the row or add another row to it

OpenStudy (anonymous):

Multiply the row by -0.25. Then you have have [0 1|4]

OpenStudy (fibonaccichick666):

that is correct! now you want any number above that 1 to be 0 can you do it?

OpenStudy (anonymous):

By just doing what we did for the second row right?

OpenStudy (fibonaccichick666):

mmhmm

OpenStudy (anonymous):

Could you just subtract the 2nd row from the first?

OpenStudy (fibonaccichick666):

yea, but you want to think of it as adding minus 1 time the second

OpenStudy (anonymous):

Yea, end up with the same answer though :P. So in the end it would be [1 0 | -19] [0 1 | 4]

OpenStudy (fibonaccichick666):

-19???

OpenStudy (anonymous):

Guess I mixed that one up too. Adding -4?

OpenStudy (fibonaccichick666):

yea but adding it to what?

OpenStudy (anonymous):

the -15...?

OpenStudy (fibonaccichick666):

ahh ok, sorry I just had no clue how you got the rest and then that so the -15 was only when we multiplied the first row by -3 that doesn't actually stay, the first row remains [1 1 5] until we add another row to it or reduce it using multiplication

OpenStudy (anonymous):

Oh. So it is only [ 1 0 | 1]? The answer would be (1,4) I assume.

OpenStudy (fibonaccichick666):

yup yup

OpenStudy (anonymous):

Thanks a bunch :). I might message you if I have a problem with another question like this. Have a nice day :)

OpenStudy (fibonaccichick666):

so this is what we are doing: http://www.purplemath.com/modules/systlin6.htm

OpenStudy (fibonaccichick666):

in case you need reference. good luck and your welcome!